solve:- (2x-7)(x-3)(x+3)(2x+5)=91

solve:-


  (2x-7)(x-3)(x+3)(2x+5)=91

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sayan

(2x-7)(x-3)(x+3)(2x+5)=91

(2x-6-1)(x-3)(x+3)(2x+6-1)=91

[2(x-3)2 - (x-3)][ 2(x+3)2 - (x+3)]=91

4(x-3)2(x+3)2  -2(x+3)2(x-3)-2(x-3)2(x+3) +(x-3)(x+3) = 91

4(x-3)2(x+3)2  -2(x+3)(x-3)2x +(x2-9) = 91

[x+2(x-3)(x+3)]2  =100

or 2x2 +x-18 =±10

 + sign

  2x2 +x-28 =0

- sign

2x2 +x-8 =0

now solve these two equation and find solution

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin